# ScaleBiO: Scalable Bilevel Optimization for LLM Data Reweighting

### v1

@@ Appendix D Proofs of Theorem 1
$\phi_{1}$ $\leq\xi_{1,k}\left(1-\frac{53\mu_{2}\eta_{0}}{72JK^{(a-c)}}\right)-\xi_{1,k}+\frac{\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}+\xi_{2,k}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}\leq 0$ (40)
$\phi_{2}$ $\leq\xi_{2,k}\left(1-\frac{17\mu_{2}\eta_{0}}{72JK^{(a-c)}}\right)-\xi_{2,k}+\frac{2\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}+\xi_{1,k}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}\leq 0$ (41)
@@
$\xi_{1,k}$ $=\frac{-\frac{53\mu_{2}\eta_{0}}{72JK^{(a-c)}}\frac{2\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}-\frac{\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}}{\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}-\frac{17\mu_{2}\eta_{0}}{72JK^{(a-c)}}\frac{53\mu_{2}\eta_{0}}{72JK^{(a-c)}}}=\frac{\frac{114\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}}{\frac{837\mu_{2}\eta_{0}}{72JK^{(a-c)}}}=\frac{10\ell^{2}\eta_{\lambda}^{0}}{\mu_{2}\eta_{0}}\frac{J}{K^{(b-a-c)}}$
$\xi_{2,k}$ $=\frac{-\frac{17\mu_{2}\eta_{0}}{72JK^{(a-c)}}\frac{\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}-\frac{2\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}}{\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}-\frac{17\mu_{2}\eta_{0}}{72JK^{(a-c)}}\frac{53\mu_{2}\eta_{0}}{72JK^{(a-c)}}}=\frac{3\ell^{2}\eta_{\lambda}^{0}}{\mu_{2}\eta_{0}}\frac{J}{K^{(b-a-c)}}$
@@
$\mathbb{E}[R_{k+1}\mid\mathcal{F}_{k},\mathcal{C}_{k}]$ $\leq R_{k}-\left(\frac{\eta_{\lambda}}{2}-\ell_{\Gamma}\eta_{\lambda}^{2}-2\xi_{1,k+1}(1+\gamma_{1})\kappa^{2}\eta_{\lambda}^{2}-2\xi_{2,k+1}(1+\gamma_{2})\kappa^{2}\eta_{\lambda}^{2}\right)\left\|\nabla\Gamma^{\alpha}(\lambda_{k})\right\|^{2}$
$+\frac{\ell_{\Gamma}\eta_{\lambda}^{2}}{2}\left(\sigma_{1}^{2}+2\alpha^{2}\sigma_{2}^{2}\right)+\xi_{1,k+1}(1+\gamma_{1}^{-1})\frac{\alpha^{2}\eta_{u}^{2}\sigma_{2}^{2}}{JB}+\xi_{2,k+1}(1+\gamma_{2}^{-1})\frac{\eta_{w}^{2}\left(\sigma_{1}^{2}+\alpha^{2}\sigma_{2}^{2}\right)}{JB}.$
@@
$\xi_{1,k}(1+\gamma_{1})\kappa^{2}\eta_{\lambda}$ $\leq\frac{40(\eta_{\lambda}^{0})^{2}K^{(a-c)}\ell^{2}\kappa^{2}J^{2}}{\mu_{2}^{2}\eta_{0}^{2}K^{b}K^{(b-a-c)}}=\frac{40(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}J^{2}}{\mu_{2}^{2}\eta_{0}^{2}K^{(2b-2a)}}\leq\frac{1}{16}$
$\xi_{2,k}(1+\gamma_{2})\kappa^{2}\eta_{\lambda}$ $\leq\frac{12(\eta_{\lambda}^{0})^{2}K^{(a-c)}\ell^{2}\kappa^{2}J^{2}}{\mu_{2}^{2}\eta_{0}^{2}K^{b}K^{(b-a-c)}}=\frac{12(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}J^{2}}{\mu_{2}^{2}\eta_{0}^{2}K^{(2b-2a)}}\leq\frac{1}{16}.$
@@
$\mathbb{E}[R_{k+1}\mid\mathcal{F}_{k},\mathcal{C}_{k}]$ $\leq R_{k}-\frac{\eta_{\lambda}}{4}\left\|\nabla\Gamma^{\alpha}(\lambda_{k})\right\|^{2}+\frac{\ell_{\Gamma}\eta_{\lambda}^{2}}{2}\left(\sigma_{1}^{2}+2\alpha^{2}\sigma_{2}^{2}\right)+\frac{8\xi_{1,k+1}}{7}\frac{\alpha^{2}\eta_{u}^{2}\sigma_{2}^{2}}{JB}+\frac{3\xi_{2,k+1}}{4}\frac{\eta_{w}^{2}\left(\sigma_{1}^{2}+\alpha^{2}\sigma_{2}^{2}\right)}{JB}.$
@@
$+\frac{4}{K\eta_{\lambda}}\sum_{k=1}^{K}\left(\frac{\ell_{\Gamma}\eta_{\lambda}^{2}}{2}\left(\sigma_{1}^{2}+2\alpha^{2}\sigma_{2}^{2}\right)+\frac{8\xi_{1,k+1}}{7}\frac{\alpha^{2}\eta_{u}^{2}\sigma_{2}^{2}}{JB}+\frac{3\xi_{2,k+1}}{4}\frac{\eta_{w}^{2}\left(\sigma_{1}^{2}+\alpha^{2}\sigma_{2}^{2}\right)}{JB}\right)$
@@
$R_{1}$ $=\Gamma^{\alpha}(\lambda_{1})-\Gamma_{\min}^{\alpha}+\xi_{1,1}\delta_{1}+\xi_{2,1}r_{1}$

### v2

@@ Appendix D Proofs of Theorem 1
$\phi_{1}$ $\leq\xi_{k}^{1}\left(1-\frac{53\mu_{2}\eta_{0}}{72JK^{(a-c)}}\right)-\xi_{k}^{1}+\frac{\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}+\xi_{k}^{2}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}\leq 0$ (39)
$\phi_{2}$ $\leq\xi_{k}^{2}\left(1-\frac{17\mu_{2}\eta_{0}}{72JK^{(a-c)}}\right)-\xi_{k}^{2}+\frac{2\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}+\xi_{k}^{1}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}\leq 0$ (40)
@@
$\xi_{k}^{1}$ $=\frac{-\frac{53\mu_{2}\eta_{0}}{72JK^{(a-c)}}\frac{2\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}-\frac{\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}}{\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}-\frac{17\mu_{2}\eta_{0}}{72JK^{(a-c)}}\frac{53\mu_{2}\eta_{0}}{72JK^{(a-c)}}}=\frac{\frac{114\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}}{\frac{837\mu_{2}\eta_{0}}{72JK^{(a-c)}}}=\frac{10\ell^{2}\eta_{\lambda}^{0}}{\mu_{2}\eta_{0}}\frac{J}{K^{(b-a-c)}}$
$\xi_{k}^{2}$ $=\frac{-\frac{17\mu_{2}\eta_{0}}{72JK^{(a-c)}}\frac{\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}-\frac{2\ell^{2}\eta_{\lambda}^{0}}{K^{(b-2c)}}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}}{\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}\frac{\mu_{2}\eta_{0}}{9JK^{(a-c)}}-\frac{17\mu_{2}\eta_{0}}{72JK^{(a-c)}}\frac{53\mu_{2}\eta_{0}}{72JK^{(a-c)}}}=\frac{3\ell^{2}\eta_{\lambda}^{0}}{\mu_{2}\eta_{0}}\frac{J}{K^{(b-a-c)}}$
@@
$\mathbb{E}[R_{k+1}\mid\mathcal{F}_{k},\mathcal{C}_{k}]$ $\leq R_{k}-\left(\frac{\eta_{\lambda}}{2}-\ell_{\Gamma}\eta_{\lambda}^{2}-2\xi_{k+1}^{1}(1+\gamma_{1})\kappa^{2}\eta_{\lambda}^{2}-2\xi_{k+1}^{2}(1+\gamma_{2})\kappa^{2}\eta_{\lambda}^{2}\right)\left\|\nabla\Gamma^{\alpha}(\lambda_{k})\right\|^{2}$
$+\frac{\ell_{\Gamma}\eta_{\lambda}^{2}}{2}\left(\sigma_{1}^{2}+2\alpha^{2}\sigma_{2}^{2}\right)+\xi_{k+1}^{1}(1+\gamma_{1}^{-1})\frac{\alpha^{2}\eta_{u}^{2}\sigma_{2}^{2}}{JB}+\xi_{k+1}^{2}(1+\gamma_{2}^{-1})\frac{\eta_{w}^{2}\left(\sigma_{1}^{2}+\alpha^{2}\sigma_{2}^{2}\right)}{JB}.$
@@
$\xi_{k}^{1}(1+\gamma_{1})\kappa^{2}\eta_{\lambda}$ $\leq\frac{40(\eta_{\lambda}^{0})^{2}K^{(a-c)}\ell^{2}\kappa^{2}J^{2}}{\mu_{2}^{2}\eta_{0}^{2}K^{b}K^{(b-a-c)}}=\frac{40(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}J^{2}}{\mu_{2}^{2}\eta_{0}^{2}K^{(2b-2a)}}\leq\frac{1}{16}$
$\xi_{k}^{2}(1+\gamma_{2})\kappa^{2}\eta_{\lambda}$ $\leq\frac{12(\eta_{\lambda}^{0})^{2}K^{(a-c)}\ell^{2}\kappa^{2}J^{2}}{\mu_{2}^{2}\eta_{0}^{2}K^{b}K^{(b-a-c)}}=\frac{12(\eta_{\lambda}^{0})^{2}\ell^{2}\kappa^{2}J^{2}}{\mu_{2}^{2}\eta_{0}^{2}K^{(2b-2a)}}\leq\frac{1}{16}.$
@@
$\mathbb{E}[R_{k+1}\mid\mathcal{F}_{k},\mathcal{C}_{k}]$ $\leq R_{k}-\frac{\eta_{\lambda}}{4}\left\|\nabla\Gamma^{\alpha}(\lambda_{k})\right\|^{2}+\frac{\ell_{\Gamma}\eta_{\lambda}^{2}}{2}\left(\sigma_{1}^{2}+2\alpha^{2}\sigma_{2}^{2}\right)+\frac{8\xi_{k+1}^{1}}{7}\frac{\alpha^{2}\eta_{u}^{2}\sigma_{2}^{2}}{JB}+\frac{3\xi_{k+1}^{2}}{4}\frac{\eta_{w}^{2}\left(\sigma_{1}^{2}+\alpha^{2}\sigma_{2}^{2}\right)}{JB}.$
@@
$+\frac{4}{K\eta_{\lambda}}\sum_{k=1}^{K}\left(\frac{\ell_{\Gamma}\eta_{\lambda}^{2}}{2}\left(\sigma_{1}^{2}+2\alpha^{2}\sigma_{2}^{2}\right)+\frac{8\xi_{k+1}^{1}}{7}\frac{\alpha^{2}\eta_{u}^{2}\sigma_{2}^{2}}{JB}+\frac{3\xi_{k+1}^{2}}{4}\frac{\eta_{w}^{2}\left(\sigma_{1}^{2}+\alpha^{2}\sigma_{2}^{2}\right)}{JB}\right)$
@@
$R_{1}$ $=\Gamma^{\alpha}(\lambda_{1})-\Gamma_{\min}^{\alpha}+\xi_{1}^{1}\delta_{1}+\xi_{2}^{1}r_{1}$